# Visual-Inertial SLAM as Simple as A, B, VINS

### v2

@@ III-D2 Optimizing the Memory Tree
where $\mathcal{S}$ is the set of necessary sequential connections to add (nodes that are next to each other temporally) and $\mathcal{L}$ is the set of loop connections. $\bm{\Sigma}_{tree}$ is the manually-tuned diagonal covariance matrix for tree edges, and $\rho_{c}$ is again the Cauchy loss function. Given a 4-DoF relative pose measurement ${}^{i}_{j}\theta_{z}$, ${}^{i}\mathbf{p}_{j}$ between nodes $i$ and $j$, the residual is
where $\{i\}$, $\{j\}$ are the local gravity-aligned frames for nodes $i$ and $j$, and $\{L\}$ is the gravity-aligned frame centered around the lowest common ancestor (LCA) for nodes $i$ and $j$. Finding the LCA can be performed in $\mathcal{O}(\log(N))$ time. Of course the poses for the nodes are defined in the parent frame and not the $\{L\}$ frame, but the poses in the $\{L\}$ frame have to be found by transforming iteratively into the same frame as the parent until the $\{L\}$ frame is reached. If optimizing all states in the memory tree, the Hessian structure will be very dense due to correlating more states than usual. Of course, the point of the memory tree is not to optimize every state, but to carefully choose which states to optimize.
Given two loop candidate tree nodes, we can actually move all the nodes needed to close the loop by only optimizing nodes along the shortest path between the two loop nodes. In order to do this, first the LCA must be found between the two. This is because the shortest path between two loop nodes in the tree always passes through the LCA. To stabilize the optimization, the root node always remains fixed. Note that this does in most cases allow the first ever pose to move, but we argue that the global pose can be simply taken relative to the first pose. Fig. 7 shows different examples of optimization patterns on the memory tree when optimizing the full path between two loop nodes. Note that edges with only fixed parameters need not be included in the optimization. Thus, not all measurements are added to the optimization problem. The path between the two loop nodes is traversed, and only at most $\mathcal{O}(\log(N))$ sequential relative edges are added, and all loop measurements are checked to see if any of their parameters are included in the set of variable parameters. If so, the loop measurement is added, and if not it is left out of the optimization problem.
While optimizing the full path between two loop nodes is already highly efficient, an even more efficient solution exists. Since optimizing a single node’s pose moves the entire subtree underneath, it is possible to start with only optimizing the LCA and its two children, and work our way down until the optimization converges in one iteration. We call this the memory tree top-down optimization algorithm. Fig. 8 shows how this works. This solution is not only highly efficient, but it also solves pose graph SLAM while only relinearizing a constant number of variables. This is because on average only three states are variable (the LCA and its two children), while the rest are fixed. However, in the worst case $\mathcal{O}(\log(N))$ states are set variable.
To robustify against potential bad loop closure measurements (e.g., from visual aliasing), we propose a robust yet efficient $\chi^{2}$ check. Before optimizing the memory tree with a new loop measurement, first the 4-DoF states of all the nodes along the shortest path between the two loop nodes are copied and stored. After optimizing the memory tree, if the new loop residual does not pass a $\chi^{2}$ test with theoretical threshold $\gamma$ (i.e., $||\mathbf{r}^{tree}_{ij}||^{2}_{\bm{\Sigma}_{tree}}\geq\gamma$ if the newest loop measurement is between nodes $i$ and $j$), then the old states are copied back into the memory tree to restore it to what it was before, and the loop measurement is discarded. Thus AB-VINS is robust to catastrophically bad loop closures while only using $\mathcal{O}(\log(N))$ extra memory.
III-D3 Analysis the Memory Tree
Table I shows the algorithm analysis of different pose graph optimization methods if optimizing with $N$ keyframes. Our memory tree top-down method is the overall best in the table. Looking at the number of variables affected by optimization (i.e., the number of variables that need to be relinearized), the standard method as in SPA [24] and iSAM [25], with the poses defined in the global frame, affects $\Theta(N)$ variables no matter what no matter what. iSAM2 [26] has a chance to affect a constant number of variables under exploration, but typically affects $\mathcal{O}(N)$ variables under a loop closure scenario. The relative method [27] can affect up to $\mathcal{O}(N)$ variables, but on average can affect only a constant number of variables due to only optimizing the beginning and end of each loop. The memory tree top-down method can also affect a constant number of variables, but in the worst case is only $\mathcal{O}(\log(N))$, and only affects a constant number of variables on average even under loop closure scenarios. Moving on to the global pose calculation, defining the poses in the global frame as is standard allows for constant-time global pose calculation. For the relative method, global pose calculation typically takes $\mathcal{O}(N)$ time since each keyframe pose is defined in the previous one’s frame of reference, and thus the global pose is dependent on each previous keyframe. On the other hand, with the memory tree, global pose calculation is typically done in $\mathcal{O}(\log(N))$ time due to the efficient tree structure of the pose definitions. It should be noted that the computation incurred by having to relinearize a larger number of variables is much higher than having to iterate to find the global pose.
The final analysis presented in Table I is the computational complexity of the overall optimization. Note that an average case is not reported here since the complexity depends on the number of loop measurements, which depends on the environment. Thus, there are two worst cases reported for $\mathcal{O}(1)$ and $\mathcal{O}(N)$ loop measurements. With a standard pose graph optimization method, the complexity is linear in the absolute best case, which is the same as having a constant number of loops. In the worst case with $\mathcal{O}(N)$ loops, the complexity is $\mathcal{O}(N^{3})$ [24]. For iSAM2 [26], the optimization can be constant-time in the best case (under exploration) but other than that shares the same complexity with the standard pose graph optimization. With the relative method, the best case arises when the two loop keyframes are very close to each other in time, and there are a constant number of loops. In this case, the method can be constant time. In the worst case with a constant number of loops, the complexity is cubic. This is because $\mathcal{O}(N)$ states can be set variable even with a single loop, and the states are all correlated leading to a dense Hessian structure. Similarly, the complexity of the relative method is cubic with $\mathcal{O}(N)$ loops. Note that this analysis of the relative method assumes that 1) it is applied to pose graph optimization and 2) only the necessary measurements are included in the optimization instead of all the measurements as proposed in the original paper [27]. If all measurements are included in the optimization then the best case becomes $\Omega(N)$ due to having to iterate through all of the sequential edges.
Moving on the the memory tree, with all states set variable, the complexity is cubic in every case. This is due of the cost of Levenberg–Marquardt iterations with a nearly dense Hessian structure. For the full path algorithm, the best case is when there are a constant number of loops and the two loop nodes are near the root of the tree. In this case, the complexity is constant. In the worst case with a constant number of loops, the complexity is $\mathcal{O}((\log(N))^{3})$, which is sublinear. This is due to the cost of Levenberg–Marquardt iterations with $\mathcal{O}(\log(N))$ states that are all correlated. In the worst case with $\mathcal{O}(N)$ loops, the complexity is $\mathcal{O}(N\log(N))$ due to having to find the LCA for $\mathcal{O}(N)$ loops to see which measurements should be included in the optimization. For the top-down method, the complexity is the same as the full path algorithm since $\mathcal{O}(\log(N))$ states can be variable in the worst case. If only a constant number of states are variable (which is the average case) with a constant number of loops, the complexity is $\mathcal{O}((\log(N))^{2})$ due to having to find the LCA for the $\mathcal{O}(\log(N))$ relative edges. While AB-VINS allows for at most $\mathcal{O}(N)$ loop measurements, some systems such as ORB-SLAM3 [2] allow for up to $\mathcal{O}(N^{2})$ loops. If this is the case, the worst case for the full path and top-down algorithms becomes $\mathcal{O}(N^{2}\log(N))$, which is still better than the cubic complexity of the other methods.
IV Experimental Results
@@
AB-VINS is implemented in an efficient C++ code base. Ceres solver [44] with single-thread execution is used for all optimizations. Automatic differentiation provided by ceres::Jet is used in all factors except for preintegration and the reprojection error from a static point used in tracking to static map, which use analytical Jacobians. Unlike numerical differentiation via finite differences, automatic differentiation does not come with a large computational overhead or loss of accuracy. We found automatic differentiation to be especially useful for developing the memory tree relative pose factor, since the analytical Jacobians for it are quite complicated. A desktop device equipped with an i5-6600K CPU and RTX 2070 Super GPU with 16GB of RAM is used in all experiments.
IV-A Accuracy
We first evaluate the accuracy of AB-VINS compared to state-of-the-art VIO [12, 43] and visual-inertial SLAM [3, 2] systems. Trajectories are aligned by finding the best 4-DoF transformation between the estimated and ground truth.
IV-A1 Pose Accuracy on the AR Table Dataset
We first test on the recently released AR table dataset [45]. This dataset has a single monocular camera with radial and tangential distortion. The results are in Table II, where it can be seen that AB-VINS is not as accurate as state-of-the-art VINS systems, either in VIO or SLAM setting. An ablation study is presented in the table where we investigate the impact of fixing each keyframe’s $\mathbf{c}$ and $\mathbf{d}$ in the local mapping window to zero separately. Clearly it is necessary to estimate both $\mathbf{c}$ and $\mathbf{d}$ to correct the monocular depth. We additionally evaluate the impact of utilizing a depth sensor instead of monocular depth, which is reported in the table. Note that $a$ and $b$ are fixed to 1 and 0 when using the depth sensor. There is a clear improvement from using more accurate depth from the depth sensor rather than the monocular depth, which shows that the biggest reason for AB-VINS being less accurate than the other systems is the reliance on monocular depth. Note that the loop-closed result of AB-VINS (using the memory tree) results in slightly improved position accuracy and slightly degraded orientation accuracy on this dataset. This shows that the memory tree did not hurt the accuracy overall in this case.
IV-A2 Pose Accuracy on the TUM-VI Dataset
The next dataset tested on is the TUM-VI dataset [46]. We use the left fisheye camera for evaluation. The results are reported in Table III, where it is clear that AB-VINS is again not as accurate as the state-of-the-art. Notice that the accuracy of AB-VINS with loop closure is significantly higher that the VIO result on this dataset. This shows the capability of the memory tree to improve the accuracy over simply running VIO. All hyperparameters of AB-VINS are the same for the TUM-VI dataset as in the AR Table dataset, which shows the generalization capability of our system.
IV-A3 Depth Accuracy
@@
The depth output of AB-VINS will ideally be accurate enough for downstream applications such as AR and path planning. We evaluate the accuracy of the depth on the AR Table dataset, which comes with ground-truth depth maps from a depth sensor. Table IV reports the results of evaluating the dense depth output of AB-VINS. We compare the base accuracy of the MiDaS DPT Swin2 Tiny model (denoted as “MiDaS” in the table) to the accuracy of the depth after jointly estimating $a$, $b$, $\mathbf{c}$, and $\mathbf{d}$ in our local mapping optimization (denoted as “AB” in the table). The scale and bias for MiDaS are estimated linearly using sparse depths from the local map (i.e., the same points as tracking to static map and depth map registration) using the linear system from the original paper [23] for a fair comparison. From Table IV it is clear that estimating the AB features results in slightly improved depth accuracy – improving all but one of the metrics in the average case. Additionally, the dense depth accuracy of both MiDaS and the AB features is amenable for downstream applications, with a $\delta_{1}$ accuracy over 0.5 and RMSE around one meter.
@@
IV-A4 Accuracy of the Memory Tree
@@
To investigate the accuracy of the memory tree, we turn to the standard 2D pose graph datasets [^4] The datasets can be downloaded here https://lucacarlone.mit.edu/datasets/. We incrementally optimize the Manhattan World (M10000 and M3500) datasets as well as the Intel and MIT Killian Court datasets to simulate large-scale SLAM settings. Optimization is only carried out when there is a new loop measurement. For the memory tree, which performs 4-DoF pose graph optimization, the $z$ positions are all set to zero in order to utilize the 2D (3-DoF) datasets. Since we only implement 4-DoF (3D position and yaw) pose graph optimization in the memory tree, the 3D (6-DoF) pose graph datasets (e.g., Sphere, Taurus, and Cube) are omitted. We compare to the state-of-the-art incremental pose graph optimization method iSAM2 [26] implemented in GTSAM [47]. We also compare the three methods for optimizing the memory tree: 1) optimizing all states, 2) optimizing the full path between two loop nodes, and 3) the top-down optimization method. The resulting trajectories for the M10000 dataset can be viewed in Fig. 9, the M3500 dataset in Fig. 10, the Intel dataset in Fig. 11, and the MIT dataset in Fig. 12. Note that the memory tree trajectories are typically tilted compared to the iSAM2 trajectory due to the fact that the first frame can not be trivially fixed and has no prior. Again, the global pose output can simply be taken relative to the first ever keyframe in order to match the output of standard methods with the first global pose fixed, but we left the trajectories tilted to show that we do not fix the first pose. Qualitatively it can be seen that optimizing all states in the memory tree results in a trajectory similar to iSAM2, while the other two memory tree methods are slightly less accurate. Of course in the Manhattan World datasets, a perfect solution would have all of the vertical parts of the trajectory at $90^{\circ}$ to the horizontal ones. However, the memory tree full path and top-down methods still produce reasonable results despite optimizing far fewer variables than is typical. On the Intel dataset, there is almost no noticeable degradation in accuracy when using the memory tree full path or top-down method, which shows a case where these methods can nearly match the state-of-the-art solution despite being far more efficient. For the MIT dataset, the performance of these two methods is noticeably worse than iSAM2 and the full batch memory tree solution. While the accuracy is still reasonable for most of the trajectory (e.g., the smaller loops), we believe that the larger loops are less accurate due to the poor odometry accuracy on this dataset. The memory tree full path and top-down methods require highly-accurate odometry in order to produce reasonable results, whereas on the MIT dataset the odemetry curves while traversing the long hallways and only optimizing a few states can not correct all of the drift needed.
IV-B Efficiency
@@
Figure 18: The trajectory plots from the static sequence of the Low/No Excitation dataset. The trajectories for AB-VINS VIO and AB-VINS are equivalent in this sequence and completely overlapping since there is only one keyframe. The trajectories look jagged only because of the extremely small scale of the plots.
@@
To evaluate efficiency, timing results are reported.
IV-B1 Efficiency of VIO Thread
@@
The main VIO thread of AB-VINS must be able to keep up with real-time requirements. Luckily, it is highly-efficient. Table V reports the efficiency of different operations of the main VIO thread – namely frame rate operations (feature tracking, vision-only, and inertial-only optimization) as well as keyframe operations (mono depth inference, depth map registration, and local mapping optimization). The total average time per frame is also reported as 11.5ms. This shows that the main VIO thread can easily keep up with real-time requirements. Note that the total average time per frame reported in Table V includes the keyframe operations (monocular depth inference, depth map registration, and local mapping optimization) in the average.
AB-VINS VIO is also efficient compared to the VIO of other state-of-the-art systems. Table VI shows the comparison. Only methods that use a single VIO thread are reported in the table, which is why ORB-SLAM3 and OKVIS are excluded. VINS-Fusion’s VIO is run in single-thread mode here to obtain the comparison. AB-VINS is much more efficient than a state-of-the-art optimization-based VINS-Fusion, and even surpasses the efficiency of a state-of-the-art filter-based OpenVINS on average while also providing dense depth. Note that launching a separate thread for local mapping as in ORB-SLAM3 would further improve the efficiency of the main thread, since only feature tracking, vision only, and inertial only would be performed on it.
IV-B2 Efficiency of the Memory Tree
@@
Where the full path and top-down memory tree methods lack in accuracy, they make up for in efficiency. Timing results can be viewed in Fig. 13 for incrementally optimizing on the 2D pose graph datasets. Here we also compare to SPA [24] for reference, which is reimplemented with Ceres solver using automatic differentiation for a fair comparison to our memory tree. SPA was omitted in the accuracy evaluation since the SPA trajectories were very similar to the iSAM2 ones. Note that SPA and iSAM2 have a slight advantage here since the implementation we use is purely 2D (3-DoF) instead of 4-DoF like ours. Optimizing all states of the memory tree is very slow – typically slower than SPA – due to the extremely large and dense Hessian structure. Both the full path and top-down methods are more efficient than SPA and iSAM2. Besides the timing, it is also interesting to look at the state size of the optimization. We claim that the top-down optimization solve pose graph SLAM while only adjusting a constant number of variables, and this is proven experimentally in Fig. 14.
@@ IV-C Robustness
IV-C1 Robustness to Degenerate Motions
@@
It is well-known that completely hand-crafted monocular VINS systems require parallax from sufficient motion to estimate the structure and poses. On the other hand, since AB-VINS estimates the structure using monocular depth, it can work even with absolutely no motion at all. To showcase this capability, we collected the Low/No Excitation dataset. The dataset has six different motion profiles. The first is static, where the device sits completely still on a table. The second is called semi-static, where the device is held as still as possible in the hand. After that, there is a sequence of rotation-only motion called rotation. Finally, there are three levels of motion ranging from 1 to 3. For motion 1, the device moves very slowly from side to side, and by motion 3 the side-to-side motion is more aggressive and includes up-and-down motion as well as rotations. All of the sequences are confined to a small area mostly facing the same thing in order to simulate a typical AR/VR scenario or low-motion robot activity (e.g., an autonomous vehicle sitting still at a traffic light or a drone hovering). The same sensor rig as the AR Table dataset is used. The results are reported in Table VII. Note that only the position error is reported here since the orientation error is impossible to calculate for the static sequences, and SE(3) alignment is used instead of the typical 4-DoF alignment since it is impossible to estimate the transform between the motion capture frame and the gravity-aligned frame with no motion in the trajectory. It can be seen in Table VII that AB-VINS (both in the VIO and SLAM setting) is the only system that can successfully initialize and estimate the poses in all sequences. A system is determined to have failed in this experiment if there was no pose output (e.g., it did not attempt to initialize) or if the position error is over 10m. OKVIS and ORB-SLAM3 fail to initialize in any of the sequences, while OpenVINS is successful in the final two, and VINS-Fusion is successful in all but the first completely static sequence and the rotation-only sequence. The performance of VINS-Fusion is surprisingly very robust to low excitation for a completely hand-crafted system, but nevertheless AB-VINS can work in all of the scenarios.
@@
We also provide visualizations of each trajectory in the dataset and the estimated trajectories for each system which successfully initialized. The plots are shown in Figs. 18, 19, 20, 21, 22, and 23 for the static, semi-static, rotation, motion 1, motion 2, and motion 3 sequences, respectively. For the static sequence in Fig. 18, AB-VINS is the only system which is able to initialize because of the lack of any motion at all. The scale of the plot is very small, which leads to a jagged appearance of the trajectories even though they are smooth from a distance. On the semi-static sequence in Fig. 19, VINS-Fusion is surprisingly able to initialize. However, it is clear from the z-axis time-series plot that it takes longer than AB-VINS to start up, since it requires at least a small amount of motion – which is present in a small amount in the z direction due to human error in trying to hold the device still. For the rotation sequence in Fig. 20, VINS-Fusion fails to initialize again. This is because the operator was able to hold the device more still in terms of translation than for the semi-static sequence, and rotation-only does not supply enough parallax for VINS-Fusion to triangulate features. On the motion 1 sequence in Fig. 21, VINS-Fusion begins to noticeable produce a better trajectory than AB-VINS, since sufficient parallax is now present. However, it again takes longer to initialize – delaying the startup of any downstream application. For the motion 2 and motion 3 sequences in Figs. 22 and 23, OpenVINS is now able to initialize and produce reasonable results. However, again, the completely hand-crafted systems exhibit delayed initialization, while AB-VINS is able to start from the first frame. Despite there being more motion in the motion 3 sequence than a typical AR/VR scenario, OKVIS and ORB-SLAM3 fail to initialize.
IV-C2 Robustness of Local Mapping Optimization

### v3

@@ V Experimental Results
V-A Accuracy
We first evaluate the localization accuracy of the proposed AB-VINS by comparing to the state-of-the-art VIO [17, 49] and visual-inertial SLAM [7, 6] methods. Trajectories are aligned by finding the best 4-DoF transformation between the estimated and ground truth, and the Absolute Trajectory Error (ATE) is used as the performance metric.
V-A1 Pose Accuracy on the AR Table Dataset
We first test on the recently released AR Table dataset [50], which has a single monocular camera with radial and tangential distortion. The results are shown in Table II, showing that our AB-VINS is not as accurate as the state-of-the-art VINS. An ablation study is presented in the table where we investigate the impact of fixing each keyframe’s $\mathbf{c}$ and $\mathbf{d}$ in the local mapping window to zero separately. Clearly it is necessary to estimate both $\mathbf{c}$ and $\mathbf{d}$ to correct the monocular depth. Additionally, we evaluate the impact of utilizing a depth sensor instead of monocular depth. Note that $a$ and $b$ are fixed to 1 and 0 when using the depth sensor. There is a clear improvement from using more accurate depth from the depth sensor rather than the monocular depth, which shows that the main reason for AB-VINS being less accurate than the other systems is the reliance on monocular depth. Interestingly, the loop closures of AB-VINS (using the memory tree) slightly improve position accuracy while degrading orientation in this case, which at least shows that the memory tree did not hurt the overall accuracy.
V-A2 Pose Accuracy on the TUM-VI Dataset
The next dataset tested on is the TUM-VI dataset [51]. We use the left fisheye camera for evaluation, and all the hyperparameters of AB-VINS are the same as in the AR Table dataset. The results are reported in Table III. Although the AB-VINS is again not as accurate as the state-of-the-art, the accuracy of AB-VINS with loop closure is significantly higher than the VIO on this dataset, showing the capability of the memory tree to improve the accuracy over simply running VIO.
V-A3 Depth Accuracy
@@
As downstream applications such as AR and path planning may require depth (which the proposed AB-VINS is able to provide), we evaluate the accuracy of the depth on the AR Table dataset. Table IV reports the results of evaluating the dense depth output of AB-VINS. We compare the base accuracy of the MiDaS DPT Swin2 Tiny model (denoted as “MiDaS” in the table) to the accuracy of the depth after jointly estimating $a$, $b$, $\mathbf{c}$, and $\mathbf{d}$ in our local mapping optimization (denoted as “AB” in the table). The scale and bias for MiDaS are estimated linearly using sparse depths from the local map (i.e., the same points as tracking to static map and depth map registration) using the linear system from the original paper [2] for a fair comparison. From Table IV it is clear that estimating the AB features results in slightly improved depth accuracy – improving all but one of the metrics in the average case. Additionally, the dense depth accuracy of both MiDaS and the AB features is amenable for many downstream applications, with a $\delta_{1}$ accuracy over 0.5 and RMSE around one meter.
@@
V-A4 Accuracy of the Memory Tree
@@
To investigate the accuracy of the memory tree, we turn to the standard 2D pose graph datasets [^4] The datasets are publicly available: https://lucacarlone.mit.edu/datasets/. We incrementally optimize the Manhattan World (M10000 and M3500) datasets as well as the Intel and MIT Killian Court datasets to simulate large-scale SLAM settings. Optimization is only carried out when there is a new loop measurement. For the memory tree, which performs 4-DoF pose graph optimization, the $z$ positions are all set to zero in order to utilize the 2D (3-DoF) datasets. Since we only implement 4-DoF (3D position and yaw) pose graph optimization in the memory tree, the 3D (6-DoF) pose graph datasets (e.g., Sphere, Taurus, and Cube) are omitted. We compare to the state-of-the-art incremental pose graph optimization method iSAM2 [31] implemented in GTSAM [52]. We also compare the three methods for optimizing the memory tree: 1) optimizing all states, 2) optimizing the full path between two loop nodes, and 3) the top-down optimization method. The resulting trajectories for the M10000 dataset can be viewed in Fig. 9, the M3500 dataset in Fig. 10, the Intel dataset in Fig. 11, and the MIT dataset in Fig. 12. Note that the memory tree trajectories are typically tilted compared to the iSAM2 trajectory due to the fact that the first frame can not be trivially fixed and has no prior. Again, the global pose output can simply be taken relative to the first ever keyframe in order to match the output of standard methods with the first global pose fixed, but we left the trajectories tilted to show that we do not fix the first pose. Qualitatively it can be seen that optimizing all states in the memory tree results in a trajectory similar to iSAM2, while the other two memory tree methods are slightly less accurate. Of course in the Manhattan World datasets, a perfect solution would have all of the vertical parts of the trajectory at $90^{\circ}$ to the horizontal ones. However, the memory tree full path and top-down methods still produce reasonable results despite optimizing far fewer variables than is typical. On the Intel dataset, there is almost no noticeable degradation in accuracy when using the memory tree full path or top-down method, which shows a case where these methods can nearly match the state-of-the-art solution despite being far more efficient. For the MIT dataset, the performance of these two methods is noticeably worse than iSAM2 and the full batch memory tree solution. While the accuracy is still reasonable for most of the trajectory (e.g., the smaller loops), we believe that the larger loops are less accurate due to the poor odometry accuracy on this dataset. The memory tree full path and top-down methods require highly-accurate odometry in order to produce reasonable results, whereas on the MIT dataset the odemetry curves while traversing the long hallways and only optimizing a few states cannot correct all of the drift needed.
V-B Efficiency
We now evaluate the efficiency of the proposed AB-VINS.
@@ V-B Efficiency
Figure 18: The trajectories from the static sequence of the Low/No Excitation dataset. The trajectories for VIO and AB-VINS are equivalent here and completely overlapping since there is only one keyframe. The trajectories look jagged only because of the extremely small scale of the plots.
@@
V-B1 Efficiency of the VIO Thread
@@
Table V reports the efficiency of different operations of the main VIO thread, namely frame rate operations (feature tracking, vision-only, and inertial-only optimization) as well as keyframe operations (mono depth inference, depth map registration, and local mapping optimization). The total average time per frame is also reported as 11.5ms, showing that the main VIO thread can easily keep up with real-time requirements. Note that the total average time per frame reported in Table V includes the keyframe operations (monocular depth inference, depth map registration, and local mapping optimization) in the average. Table VI shows the comparison against the state-of-the-art VIO systems. Only methods that use a single VIO thread are reported in the table, which is why ORB-SLAM3 and OKVIS are excluded. VINS-Fusion’s VIO is run in single-thread mode here to obtain the comparison. AB-VINS is much more efficient than a state-of-the-art optimization-based VINS-Fusion, and even surpasses the efficiency of a state-of-the-art filter-based OpenVINS on average while also providing dense depth. Note that launching a separate thread for local mapping as in ORB-SLAM3 would further improve the efficiency of the main thread, since only feature tracking, vision only, and inertial only would be performed on it.
V-B2 Efficiency of the Memory Tree
@@
Where the full path and top-down memory tree methods lack in accuracy, they make up for in efficiency. Timing results can be viewed in Fig. 13 for incrementally optimizing on the 2D pose graph datasets. Here we also compare to SPA [29] for reference, which is reimplemented with Ceres solver using automatic differentiation for a fair comparison to our memory tree. SPA was omitted in the accuracy evaluation since the SPA trajectories were very similar to the iSAM2 ones. Note that SPA and iSAM2 have a slight advantage here since the implementation we use is purely 2D (3-DoF) instead of 4-DoF like ours. Optimizing all states of the memory tree is very slow – typically slower than SPA – due to the extremely large and dense Hessian structure. Both the full path and top-down methods are more efficient than SPA and iSAM2. Besides the timing, it is also interesting to look at the state size of the optimization. We claim that the top-down optimization solves pose graph SLAM while only adjusting a constant number of variables, and this is proven experimentally in Fig. 14.
@@ V-C Robustness
V-C1 Robustness to Degenerate Motions
@@
It is known that completely hand-crafted monocular VINS require parallax from sufficient motion to estimate the structure and poses. On the other hand, since AB-VINS estimates the structure using monocular depth, it can work even with absolutely no motion at all. To showcase this capability, we collected the Low/No Excitation dataset. The dataset has six different motion profiles. The first is static, where the device sits completely still on a table. The second is called semi-static, where the device is held as still as possible in the hand. After that, there is a sequence of rotation-only motion called rotation. Finally, there are three levels of motion ranging from 1 to 3. For motion 1, the device moves very slowly from side to side, and by motion 3 the side-to-side motion is more aggressive and includes up-and-down motion as well as rotations. All of the sequences are confined to a small area mostly facing the same thing in order to simulate a typical AR/VR scenario or low-motion robot activity (e.g., an autonomous vehicle sitting still at a traffic light or a drone hovering). The same sensor rig as the AR Table dataset is used. The results are reported in Table VII. Note that only the position error is reported here since the orientation error is impossible to calculate for the static sequences, and SE(3) alignment is used instead of the typical 4-DoF alignment since it is impossible to estimate the transform between the motion capture frame and the gravity-aligned frame with no motion in the trajectory. It can be seen in Table VII that AB-VINS (both in the VIO and SLAM setting) is the only system that can successfully initialize and estimate the poses in all sequences. A system is determined to have failed in this experiment if there was no pose output (e.g., it did not attempt to initialize) or if the position error is over 10m. OKVIS and ORB-SLAM3 fail to initialize in any of the sequences, while OpenVINS is successful in the final two, and VINS-Fusion is successful in all but the first completely static sequence and the rotation-only sequence. The performance of VINS-Fusion is surprisingly very robust to low excitation for a completely hand-crafted system, but nevertheless AB-VINS can work in all of the scenarios.
@@
We also provide visualizations of each trajectory in the dataset and the estimated trajectories for each system which successfully initialized. The plots are shown in Figs. 18, 19, 20, 21, 22, and 23 for the static, semi-static, rotation, motion 1, motion 2, and motion 3 sequences, respectively. For the static sequence in Fig. 18, AB-VINS is the only system which is able to initialize because of the lack of any motion at all. The scale of the plot is very small, which leads to a jagged appearance of the trajectories even though they are smooth from a distance. On the semi-static sequence in Fig. 19, VINS-Fusion is surprisingly able to initialize. However, it is clear from the $z$-axis time-series plot that it takes longer than AB-VINS to start up, since it requires at least a small amount of motion – which is present in a small amount in the $z$ direction due to human error in trying to hold the device still. For the rotation sequence in Fig. 20, VINS-Fusion fails to initialize again. This is because the operator was able to hold the device more still in terms of translation than for the semi-static sequence, and rotation-only does not supply enough parallax for VINS-Fusion to triangulate features. On the motion 1 sequence in Fig. 21, VINS-Fusion begins to noticeably produce a better trajectory than AB-VINS, since sufficient parallax is now present. However, it again takes longer to initialize – delaying the startup of any downstream application. For the motion 2 and motion 3 sequences in Figs. 22 and 23, OpenVINS is now able to initialize and produce reasonable results. However, again, the completely hand-crafted systems exhibit delayed initialization, while AB-VINS is able to start from the first frame. Despite there being more motion in the motion 3 sequence than a typical AR/VR scenario, OKVIS and ORB-SLAM3 fail to initialize.
V-C2 Robustness of Local Mapping Optimization